OpenStudy (anonymous):

if 3x+2y=11 and 2x+3y=17, what is the average (arithmetic mean) of x and y?

OpenStudy (watchmath):

Add the two equations together \(5x+5y=28\) Hence the average mean \((x+y)/2=28/10=14/5\)

OpenStudy (anonymous):

yeah